solve for x
x^x^x^x^.....x^3 = 3

Answer by Nate(3500)

x^x^x^x^.....x^3 = 3
log( x^x^x^x^.....x^3 ) = log( 3 )
x^x^x^.....x^3 * log( x ) = log( 3 )
3 * log( x ) = log( 3 )
log( x ) = log( 3 )/3
x = 10^( log( 3 )/3 )
About 1.4422